# Visualizing Quantum States: A Pilot Study on Problem Solving in Quantum Information Science Education

### v2

@@ A.1 One-qubit survey
| survey part | A | B | combined | |||
| Global Phase Equiv. | 0.56 | 0.5 | 0.62 | 1.0 | 0.59 | 0.76 |
| X gate | 0.56 | 0.25 | - | - | - | - |
| Z gate | - | - | 0.62 | 0.78 | - | - |
| phase gate | 0.78 | 0.75 | 0.88 | 0.78 | 0.82 | 0.76 |
| H gate | 0.67 | 0.62 | 0.75 | 0.78 | 0.71 | 0.71 |
| Measurement Prob. | 0.56 | 0.12 | 0.38 | 0.67 | 0.47 | 0.41 |
| Measurement State | 0.89 | 0.75 | 0.75 | 1.0 | 0.82 | 0.88 |
| Mean | 0.670.13 | 0.50.24 | 0.670.16 | 0.830.12 | 0.680.15 | 0.710.18 |
| survey part | A | B | combined | |||
@@
A.1.1 Representational competence
Participants were asked to calculate Euler formula identities and similar questions to translate between CN to a complex number. In addition, they were asked to multiply numbers written in DN and CN as shown in Figure 7. Then they were asked to rate their confidence in the question from 0 (random guess), 1 (very unsure) to 4 (very sure). The correct answers were weighted (multiplied by , where is the confidence). The weighted (unweighted) average mathematics representational understanding was (). The weighted (unweighted) average CN representational understanding was (). The average weighted (unweighted) difference of CN representational understanding minus mathematics representational understanding was ().
@@ A.1.1 Representational competence
A.1.2 Translation competence and cognitive load
On average translation competence, 12 students achieved perfect scores and 5 students achieved a score of 0.5. The students who did not score perfectly also scored lower in the rest of the survey ( to without visualization and to with visualization) and took a smaller amount of time on the correctly solved questions ( to seconds without visualization and to seconds with visualization) than the students who solved the translation questions perfectly.
A.2 Two-qubit survey
@@
| survey part | A | B | combined | |||
| X gate | 1.0 | 0.4 | 0.6 | 0.8 | 0.8 | 0.6 |
| H gate | 0.8 | 0.2 | 0.4 | 0.6 | 0.6 | 0.4 |
| Measurement Prob. | 0.6 | 0.6 | 0.4 | 0.6 | 0.5 | 0.6 |
| Separability | 0.6 | 1.0 | 0.8 | 0.6 | 0.7 | 0.8 |
| Entanglement | 0.8 | 1.0 | 0.6 | 0.4 | 0.7 | 0.7 |
| CNOT gate | 0.8 | 0.8 | 0.8 | 0.6 | 0.8 | 0.7 |
| Guess-the-Gates | 0.6 | 0.6 | 0.4 | 0.6 | 0.5 | 0.6 |
| Mean | 0.740.14 | 0.660.28 | 0.570.17 | 0.60.11 | 0.660.12 | 0.630.12 |
@@
| survey part | A | B | combined | |||
| question | w/o vis.333N = 6 | with vis.222N = 5 | w/o vis.222N = 5 | with vis.333N = 6 | w/o vis.444Standard deviation is used for confidence intervals.Confidence intervals are the average of the standard deviations. | with vis.444N = 3 |
| X gate | 1.0 | 0.8 | 0.8 | 0.83 | 0.91 | 0.82 |
| H gate | 0.17 | 0.6 | 0.4 | 0.17 | 0.27 | 0.36 |
| Measurement Prob. | 0.5 | 0.6 | 0.6 | 0.33 | 0.55 | 0.45 |
| Separability | 0.5 | 0.4 | 0.6 | 0.83 | 0.55 | 0.64 |
| Entanglement | 0.5 | 0.8 | 0.6 | 0.33 | 0.55 | 0.55 |
| CNOT gate | 0.67 | 1.0 | 0.8 | 0.5 | 0.73 | 0.73 |
| Guess-the-Gates | 0.67 | 0.2 | 0.0 | 0.17 | 0.36 | 0.18 |
| Mean | 0.570.23 | 0.630.25 | 0.540.26 | 0.450.26 | 0.560.2 | 0.530.2 |
@@
| survey part | A | B | combined | |||
| question | w/o vis.222N = 6 | with vis.333N = 11 | w/o vis.333N = 5 | with vis.222N = 6 | w/o vis.111N = 11 | with vis.111N = 11 |
@@
| survey part | A | B | combined | |||
@@ A.2.1 Circle Notation
A.2.2 Dimensional Circle Notation
@@
A.3 Three-qubit survey
@@
| survey part | A | B | combined | |||
@@
| Dimensional Circle Notation222N = 5 | ||||||
| survey part | A | B | combined | |||
| X gate | 1.0 | 0.67 | 0.67 | 1.0 | 0.8 | 0.8 |
| Z gate | 1.0 | 0.67 | 0.33 | 1.0 | 0.6 | 0.8 |
| H gate | 1.0 | 0.0 | 0.67 | 0.5 | 0.8 | 0.2 |
| Measurement Prob. | 1.0 | 0.67 | 0.67 | 1.0 | 0.8 | 0.8 |
| Measurement State | 0.5 | 1.0 | 0.33 | 1.0 | 0.4 | 1.0 |
| Separability | 1.0 | 0.33 | 0.67 | 0.0 | 0.8 | 0.2 |
| Entanglement | 1.0 | 0.67 | 0.67 | 1.0 | 0.8 | 0.8 |
| CNOT gate | 1.0 | 0.67 | 0.33 | 0.5 | 0.6 | 0.6 |
| Guess-the-Gates | 0.5 | 0.33 | 0.67 | 0.5 | 0.6 | 0.4 |
| Mean | 0.890.21 | 0.560.27 | 0.560.16 | 0.720.34 | 0.690.15 | 0.620.29 |
@@
| survey part | A | B | combined | |||
@@ A.3.1 Circle Notation
Group vis-math, that was presented questions with CN first, scored higher on the H gate, CNOT gate, and Guess-the-Gates questions with CN with to . The group scored lower with CN on the Measurement State question with 0.6 as compared to 0.8. When presented with CN, participants in the group took longer on the Z gate, H gate, Measurement Prob., Measurement State, Separability, and CNOT gate questions, with to . They took less time on the X gate, entanglement, and Guess-the-Gates questions, with to seconds.
A.3.2 Dimensional Circle Notation

### v3

@@ B.1 One-qubit survey
| survey part: | A | B | combined | |||
| question | w/o vis.111N = 12 | with vis.222N = 11 | w/o vis.222N = 11 | with vis.111N = 12 | w/o vis.333N = 23 | with vis.333N = 23 |
| Global Phase | 0.62 | 0.67 | 0.57 | 1.0 | 0.59 | 0.84 |
| X gate | 0.46 | 0.29 | - | - | - | - |
| Z gate | - | - | 0.54 | 0.75 | - | - |
| Phase gate | 0.77 | 0.64 | 0.92 | 0.83 | 0.85 | 0.73 |
| H gate | 0.69 | 0.57 | 0.54 | 0.67 | 0.62 | 0.62 |
| Measurement Prob. | 0.62 | 0.21 | 0.5 | 0.67 | 0.56 | 0.42 |
| Measurement State | 0.92 | 0.71 | 0.64 | 1.0 | 0.78 | 0.85 |
| Mean | 0.680.14 | 0.520.19 | 0.620.14 | 0.820.14 | 0.680.11 | 0.690.16 |
| survey part: | A | B | combined | |||
@@
B.1.1 Representational competence
Participants were asked to calculate Euler formula identities and similar questions to translate between CN to a complex number. In addition, they were asked to multiply numbers written in DN and CN as shown in Figure 7. Then they were asked to rate their confidence in the question from 0 (random guess), 1 (very unsure) to 4 (very sure). The correct answers were weighted (multiplied by , where is the confidence). The weighted (unweighted) average symbolic representational understanding was (). The weighted (unweighted) average CN representational understanding was (). The average weighted (unweighted) difference of CN representational understanding minus mathematics representational understanding was ().
@@ B.1.1 Representational competence
B.1.2 Connectional understanding and cognitive load
On average connectional understanding, 12 students achieved perfect scores and 5 students achieved a score of 0.5. The students who did not score perfectly also scored lower in the rest of the survey ( to without visualization and to with visualization) and took a smaller amount of time on the correctly solved questions ( to seconds without visualization and to seconds with visualization) than the students who solved the connectional understanding questions perfectly.
B.2 Two-qubit survey
@@
| survey part: | A | B | combined | |||
| question | w/o vis.222N = 6 | with vis.222N = 6 | w/o vis.222N = 6 | with vis.222N = 6 | w/o vis.111N = 12 | with vis.111N = 12 |
| X gate | 0.86 | 0.5 | 0.67 | 0.86 | 0.77 | 0.69 |
| H gate | 0.86 | 0.17 | 0.33 | 0.57 | 0.62 | 0.38 |
| Measurement Prob. | 0.57 | 0.5 | 0.33 | 0.57 | 0.46 | 0.54 |
| Separability | 0.57 | 1.0 | 0.83 | 0.57 | 0.69 | 0.77 |
| Entanglement | 0.57 | 0.83 | 0.5 | 0.29 | 0.54 | 0.54 |
| Guess-the-Gates | 0.5 | 0.5 | 0.33 | 0.5 | 0.42 | 0.5 |
| Mean | 0.650.15 | 0.580.27 | 0.50.19 | 0.560.17 | 0.580.12 | 0.570.13 |
@@
| survey part: | A | B | combined | |||
| question | w/o vis.444N = 8 | with vis.333N = 7 | w/o vis.333N = 7 | with vis.444N = 8 | w/o vis.555N = 15 | with vis.555N = 15 |
| X gate | 1.0 | 0.88 | 0.88 | 0.71 | 0.93 | 0.8 |
| H gate | 0.14 | 0.62 | 0.5 | 0.14 | 0.33 | 0.4 |
| Measurement Prob. | 0.43 | 0.5 | 0.62 | 0.29 | 0.53 | 0.4 |
| Separability | 0.57 | 0.62 | 0.75 | 0.71 | 0.67 | 0.67 |
| Entanglement | 0.43 | 0.88 | 0.75 | 0.43 | 0.6 | 0.67 |
| Guess-the-Gates | 0.71 | 0.2 | 0.0 | 0.17 | 0.42 | 0.18 |
| Mean | 0.550.27 | 0.620.23 | 0.580.29 | 0.410.24 | 0.580.19 | 0.520.21 |
@@
| survey part: | A | B | combined | |||
@@
| survey part: | A | B | combined | |||
@@ B.2.1 Circle Notation
B.2.2 Dimensional Circle Notation
@@
B.3 Three-qubit survey
@@
| survey part: | A | B | combined | |||
| question | w/o vis.222N = 7 | with vis.222N = 7 | w/o vis.222N = 7 | with vis.222N = 7 | w/o vis.111N = 10 | with vis.111N = 10 |
@@
| Dimensional Circle Notation222N = 7 | ||||||
| survey part: | A | B | combined | |||
| question | w/o vis.333N = 3 | with vis.444N = 4 | w/o vis.444N = 4 | with vis.333N = 3 | w/o vis.222N = 7 | with vis.222N = 7 |
| X gate | 1.0 | 0.75 | 0.75 | 1.0 | 0.86 | 0.86 |
| Z gate | 1.0 | 0.75 | 0.5 | 0.67 | 0.71 | 0.71 |
| H gate | 1.0 | 0.25 | 0.75 | 0.67 | 0.86 | 0.43 |
| Measurement Prob. | 1.0 | 0.75 | 0.75 | 1.0 | 0.86 | 0.86 |
| Measurement State | 0.33 | 1.0 | 0.5 | 1.0 | 0.43 | 1.0 |
| Separability | 1.0 | 0.5 | 0.75 | 0.0 | 0.86 | 0.29 |
| Entanglement | 0.67 | 0.75 | 0.75 | 1.0 | 0.71 | 0.86 |
| CNOT gate | 1.0 | 0.75 | 0.5 | 0.67 | 0.71 | 0.71 |
| Guess-the-Gates | 0.33 | 0.25 | 0.75 | 0.33 | 0.57 | 0.29 |
| Mean | 0.810.28 | 0.640.24 | 0.670.12 | 0.70.33 | 0.730.14 | 0.670.25 |
@@
| survey part: | A | B | combined | |||
@@ B.3.1 Circle Notation
Group vis-math, that was presented questions with CN first, scored higher on the H gate, CNOT gate, and Guess-the-Gates questions with CN with to . The group scored lower with CN on the Measurement State question with 0.6 as compared to 0.8. When presented with CN, participants in the group took longer on the Z gate, H gate, Measurement Prob., Measurement State, Separability, and CNOT gate questions, with to seconds. They took less time on the X gate, entanglement, and Guess-the-Gates questions, with to seconds.
B.3.2 Dimensional Circle Notation